Gaming
Probability System Calculator
Calculate gacha rates, drop chances, and success probabilities in games
Expected Attempts2.5 attemptsView results
%
Real-time Results
Expected Attempts
2.5attempts
Average attempts needed for target successes
Expected Total Cost
2,500gold/currency
Failure Penalty
Cumulative Success Probability
Probability of at least 1 success by attempt N
| Item | Value |
|---|---|
| Prob after 1 attempts | 40.00% |
| Prob after 2 attempts | 64.00% |
| Prob after 3 attempts | 78.40% |
| Prob after 5 attempts | 92.22% |
| Prob after 10 attempts | 99.40% |
| Prob after 20 attempts | 100.00% |
| Prob after 50 attempts | 100.00% |
| Prob after 100 attempts | 100.00% |
Supported inputs
| Input | Description |
|---|---|
| Success Rate per Attempt (%) (%) | Probability of a single enhancement attempt succeeding |
| Cost per Attempt | Gold / currency per enhancement attempt |
| Target Successes | Enter the number of successful attempts you want to achieve |
| Failure Penalty | The consequence or penalty when an enhancement attempt fails nothing · downgrade · destroy |
Formula
Expected attempts = 1 / p Expected total attempts = (1/p) × successes needed Total cost = Expected attempts × cost per attempt P(at least 1 in n attempts) = 1 − (1−p)ⁿ
Related Calculators
More calculators selected by category and topic.
Related Tags:
